OpenStudy (anonymous):

find the total area of the cylinder given the radius(r) or diameter(d) and height (h) r=5ft h=12ft

OpenStudy (bahrom7893):

area of a cylinder is the sum of top and bottom circles and a folded rectangle that form it

OpenStudy (bahrom7893):

A of one circle: pi*r^2=25pi A of the retangle: 2pi*r*h=120pi

OpenStudy (bahrom7893):

So total area is: 2*25pi+120pi=170pi